# Learn2Mix: Training Neural Networks Using Adaptive Data Integration

### v2

@@ 10            $\smash{S^{p}\leftarrow\biguplus_{i=1}^{k}S_{i}^{p}}$   (Aggregate to form batch $S^{p}$)
13       $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta,\alpha)$   (Update model params, $\theta$)
@@
15             $\mathcal{L}_{i}(\theta)\leftarrow\frac{1}{P}\sum_{p=1}^{P}\frac{1}{\alpha_{i}M}\sum_{(x_{j},y_{j})\in S_{i}^{p}}\ell(h_{\theta}(x_{j}),y_{j})$   (Compute loss for class $i$)
@@ 16      $\alpha\leftarrow\texttt{Update\_Mixing\_Params}(\alpha,\mathcal{L}(\theta),\gamma)$
Algorithm 2 encapsulates the mechanism for adjusting class proportions via the mixing parameters, $\alpha$, based on the computed class-wise losses. For each class, $i\in\{1,\ldots,k\}$, the algorithm normalizes the class-wise loss, $\mathcal{L}_{i}(\theta)$, by the cumulative loss across classes to obtain $L_{i}$. The mixing parameter $\alpha_{i}$ is then updated by moving it towards $L_{i}$, with the step size controlled by the mixing rate, $\gamma$. This adaptive update ensures that classes with higher error rates receive increased attention in subsequent epochs, promoting balanced and focused learning across all classes.
@@
This cyclic selection procedure ensures that the required number of samples, $\alpha_{i}M$, for each class in every batch is maintained, even as $\alpha_{i}$ is adapyively updated across epochs.
@@
3       $\alpha_{i}\leftarrow\alpha_{i}+\gamma\left(L_{i}-\alpha_{i}\right)$   (Update mixing params)
@@
Algorithm 2 Updating Mixing Params Via Learn2Mix
@@ 4 Empirical Results
In this section, we present our empirical results on classification, regression, and image reconstruction tasks, across both benchmark and modified class imbalanced datasets. We first present the classification results on three benchmark datasets (MNIST [Deng, 2012], Fashion-MNIST [Xiao et al., 2017], CIFAR-10 [Krizhevsky et al., 2009]), and three standard datasets with manually imbalanced classes (Imagenette [Howard, 2020], CIFAR-100 [Krizhevsky et al., 2009], and IMDB [Maas et al., 2011]). We note that for the imbalanced case, we only introduce the manual class-imbalancing to the training dataset, $J$, where the test dataset, $\smash{K=\{(x_{j},y_{j})\}_{j=1}^{N_{\text{test}}}}$, is not changed. This choice ensures that the generalization performance of the network is benchmarked in a class-balanced setting. Next, for the regression task, we study two benchmark datasets with manually imbalanced classes (Wine Quality [Cortez et al., 2009], and California Housing [Géron, 2022]), and a synthetic mean estimation task, wherein the manual class-imbalancing parallels that of the classification case. Finally, we reconsider the MNIST, Fashion MNIST and CIFAR-10 datasets in the context of image reconstruction, again considering the aforementioned manual class-imbalancing procedure. A comprehensive description of these datasets and class-imbalancing strategies is provided in Section B of the Appendix.
@@
The intuition behind the application of learn2mix to regression and reconstruction tasks stems from its ability to adaptively handle different data distributions. As an example, for regression tasks involving a categorical variable taking $k$ distinct values, the samples from the original training dataset, $J$, that correspond to each of these $k$ values, can be aggregated to obtain each class-specific training dataset, $J_{i}$. Here, each dataset, $J_{i}$, represents a different underlying distribution. Paralleling the classification case, learn2mix will adaptively adjust the proportions of the class-specific datasets during training. Similarly, in the context of image reconstruction, we can treat the $k$ distinct classes being reconstructed as the values taken by a categorical variable, paralleling the regression context. This formulation supports the adaptive adjustment of class proportions under learn2mix training.
@@
For the evaluations that follow, to ensure a fair comparison between the learn2mix training strategy and the classical training strategy, we use the same learning rate, $\eta$, and neural network architecture with initialized parameters, $\theta$, across all experiments for a given dataset. Additionally, we train each neural network through learn2mix (with mixing rate $\gamma$) and classical training for $E$ training epochs, where $E$ is dataset and task dependent [^1] Practically, we observe that choosing $\gamma\in[0.01,0.5]$ yields improved performance (see empirical results).. In classification tasks, we also benchmark learn2mix and classical training versus ‘focal training’ and ‘SMOTE training’ (training using focal loss [Lin et al., 2017] and SMOTE oversampling [Chawla et al., 2002] — see Sections C.2 and C.3 of the Appendix). The complete list of model architectures and hyperparameter choices is provided in Section C of the Appendix.
@@
Table 1: Test classification accuracy for learn2mix (L2M), classical (CL), focal (FCL), SMOTE training.
Epoch $\boldsymbol{t=0.25E}$
Dataset MNIST Fashion MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Accuracy (L2M) $77.62\scriptstyle\pm 1.83$ $46.52\scriptstyle\pm 3.25$ $51.38\scriptstyle\pm 0.40$ $33.89\scriptstyle\pm 1.66$ $7.270\scriptstyle\pm 0.46$ $70.82\scriptstyle\pm 1.69$
Accuracy (CL) $66.07\scriptstyle\pm 4.57$ $40.54\scriptstyle\pm 3.43$ $49.89\scriptstyle\pm 0.51$ $25.16\scriptstyle\pm 1.01$ $4.600\scriptstyle\pm 0.32$ $53.82\scriptstyle\pm 3.93$
Accuracy (FCL) $69.92\scriptstyle\pm 4.71$ $40.59\scriptstyle\pm 3.42$ $49.59\scriptstyle\pm 0.70$ $27.63\scriptstyle\pm 2.15$ $6.836\scriptstyle\pm 0.27$ $50.89\scriptstyle\pm 1.10$
Accuracy (SMOTE) $67.87\scriptstyle\pm 5.23$ $40.43\scriptstyle\pm 3.47$ $50.08\scriptstyle\pm 0.53$ $29.76\scriptstyle\pm 0.72$ $6.570\scriptstyle\pm 0.42$ $54.38\scriptstyle\pm 2.41$
Epoch $\boldsymbol{t=0.5E}$
Dataset MNIST Fashion MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Accuracy (L2M) $85.04\scriptstyle\pm 1.38$ $60.12\scriptstyle\pm 1.30$ $56.76\scriptstyle\pm 0.69$ $43.50\scriptstyle\pm 0.86$ $12.10\scriptstyle\pm 0.36$ $76.12\scriptstyle\pm 2.36$
Accuracy (CL) $82.69\scriptstyle\pm 1.58$ $54.59\scriptstyle\pm 3.11$ $55.36\scriptstyle\pm 0.40$ $33.72\scriptstyle\pm 1.24$ $8.200\scriptstyle\pm 0.26$ $72.32\scriptstyle\pm 3.28$
Accuracy (FCL) $83.46\scriptstyle\pm 1.52$ $56.09\scriptstyle\pm 2.56$ $54.81\scriptstyle\pm 0.43$ $35.82\scriptstyle\pm 0.97$ $11.12\scriptstyle\pm 0.40$ $69.33\scriptstyle\pm 3.89$
Accuracy (SMOTE) $82.93\scriptstyle\pm 1.67$ $54.55\scriptstyle\pm 3.10$ $54.76\scriptstyle\pm 0.68$ $38.73\scriptstyle\pm 0.47$ $10.86\scriptstyle\pm 0.47$ $66.28\scriptstyle\pm 1.78$
Epoch $\boldsymbol{t=E}$
Dataset MNIST Fashion MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Accuracy (L2M) $91.18\scriptstyle\pm 1.03$ $67.34\scriptstyle\pm 1.18$ $62.10\scriptstyle\pm 0.39$ $53.31\scriptstyle\pm 0.68$ $17.02\scriptstyle\pm 0.48$ $82.33\scriptstyle\pm 0.50$
Accuracy (CL) $90.01\scriptstyle\pm 1.12$ $65.27\scriptstyle\pm 1.74$ $61.46\scriptstyle\pm 0.31$ $44.60\scriptstyle\pm 0.68$ $12.62\scriptstyle\pm 0.37$ $80.03\scriptstyle\pm 0.48$
Accuracy (FCL) $90.08\scriptstyle\pm 1.07$ $66.32\scriptstyle\pm 1.71$ $61.19\scriptstyle\pm 0.18$ $45.30\scriptstyle\pm 0.74$ $14.45\scriptstyle\pm 0.57$ $79.83\scriptstyle\pm 0.71$
Accuracy (SMOTE) $90.08\scriptstyle\pm 1.13$ $65.27\scriptstyle\pm 1.73$ $60.93\scriptstyle\pm 0.25$ $49.41\scriptstyle\pm 0.73$ $15.05\scriptstyle\pm 0.61$ $77.46\scriptstyle\pm 0.70$
@@
Figure 2: Comparing model classification accuracies across six datasets (MNIST, Fashion MNIST, CIFAR-10, Imagenette, CIFAR-100, IMDB Sentiment Analysis) using Cross Entropy Loss for classical training, learn2mix training, focal training, and SMOTE training. The x-axis is the number of elapsed training epochs, while the y-axis is the classification accuracy.
@@ 4.1 Classification Tasks
As illustrated in Table 1 and Figure 2, we observe a consistent trend across all considered classification benchmarks, whereby neural networks trained using learn2mix converge faster than their classically-trained, focal loss-trained, and SMOTE-trained counterparts. More concretely, we begin by considering the MNIST benchmark dataset. We train LeNet-5 [Lecun et al., 1998] via the Adam optimizer [Kingma and Ba, 2014] and Cross Entropy Loss for $E=60$ epochs on MNIST, leveraging learn2mix, classical, focal, and SMOTE training. We note that the learn2mix-trained CNN achieves faster convergence, eclipsing a test accuracy of $75\%$ after 14 epochs, whereas the respective classically-trained, focal loss-trained, and SMOTE-trained CNNs achieve this test accuracy after 20 epochs, 18 epochs, and 19 epochs. Subsequently, we consider the more challenging Fashion MNIST benchmark dataset. We train LeNet-5 for $E=70$ epochs with the Adam optimizer and Cross Entropy Loss on Fashion MNIST, leveraging learn2mix, classical, focal, and SMOTE training. Paralleling the MNIST case, we observe that the learn2mix-trained CNN achieves faster convergence, yielding a test accuracy of $60\%$ after 35 epochs, whereas the respective classically-trained, focal loss-trained, and SMOTE-trained CNNs achieve this test accuracy after 49 epochs, 44 epochs, and 49 epochs. The last class-balanced benchmark dataset we investigate is the CIFAR-10 dataset, which offers a greater challenge than MNIST and Fashion MNIST. We train LeNet-5 for $E=125$ epochs using the Adam optimizer and Cross Entropy Loss on CIFAR-10, utilizing learn2mix, classical, focal, and SMOTE training. We observe that the learn2mix-trained CNN achieves faster convergence, yielding a test accuracy of $55\%$ after 50 epochs, whereas the respective classically-trained, focal loss-trained, and SMOTE-trained CNNs exceed this test accuracy after 60 epochs, 61 epochs, and 60 epochs. Cumulatively, these evaluations demonstrate the efficacy of learn2mix training in settings with balanced classes, wherein the adaptive adjustment of class proportions accelerates convergence.
@@
We now consider the case of benchmarking classification accuracies when the training dataset consists of imbalanced classes. We first benchmark the Imagenette dataset, which is characterized by a subset of 10 classes from ImageNet [Deng et al., 2009], and modify the training dataset such that the number of samples from each class, $i\in\{1,\ldots,k\}$, in $J$ decreases linearly. We train ResNet-18 [He et al., 2016] utilizing the Adam optimizer and Cross Entropy Loss for $E=60$ epochs on Imagenette, via learn2mix, classical, focal, and SMOTE training. We observe that the learn2mix-trained ResNet-18 model converges faster, achieving a test accuracy of $40\%$ after 22 epochs, at which point the respective classically-trained, focal loss-trained, and SMOTE-trained ResNet-18 models have test accuracies of $30\%$, $32\%$ and $35\%$. Next, we consider the CIFAR-100 dataset, and again modify the training dataset such that the number of samples from each class, $i\in\{1,\ldots,k\}$, in $J$ decreases logarithmically. We train LeNet-5 for $E=120$ epochs using the Adam optimizer and Cross Entropy Loss on CIFAR-100, via learn2mix, classical, focal, and SMOTE training. We see that the learn2mix-trained LeNet-5 model observes faster convergence, achieving a test accuracy of $15\%$ after 90 epochs, at which point the respective classically-trained, focal loss-trained, and SMOTE-trained CNNs have test accuracies of $11\%$ and $13.3\%$, and $13.4\%$. We further note that the $k=100$ mixing parameters within learn2mix are a small fraction of the total model parameters, making this overhead negligible. Regarding the IMDB dataset, we modify the training dataset such that the positive class keeps $30\%$ of its original samples. We train a transformer for $E=40$ epochs utilizing the Adam optimizer and Cross Entropy Loss on IMDB, with learn2mix, classical, focal, and SMOTE training. We find that the learn2mix-trained transformer converges faster, reaching a test accuracy of $75\%$ after 16 epochs, at which point the respective classically-trained, focal loss-trained, and SMOTE-trained transformers have test accuracies of $68\%$, $62\%$, and $61.8\%$. These experiments demonstrate the efficacy of learn2mix training over existing approaches in imbalanced classification settings.
@@
In the class-imbalance evaluations, we see that learn2mix not only accelerates convergence, but also achieves a tighter alignment between training and test errors compared to classical training. This correspondence indicates reduced overfitting, as learn2mix inherently adjusts class proportions based on class-specific error rates, $L_{i}$. By biasing the optimization procedure away from the original class distribution and towards $L_{i}$, learn2mix improves the model’s generalization performance. We note that this property is not unique to classification and also applies to regression and reconstruction (see Sections 4.2 and 4.3).

### v3

@@ 10    $\smash{S^{p}\leftarrow\biguplus_{i=1}^{k}S_{i}^{p}}$ (Aggregate to form batch $S^{p}$)
13 $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta,\alpha)$ (Update model parameters, $\theta$)
@@
15    $\mathcal{L}_{i}(\theta)\leftarrow\frac{1}{P}\sum_{p=1}^{P}\frac{1}{\alpha_{i}M}\sum_{(x_{j},y_{j})\in S_{i}^{p}}\ell(h_{\theta}(x_{j}),y_{j})$ (Compute loss for class ii)
@@ 16 $\alpha\leftarrow\texttt{Update\_Mixing\_Params}(\alpha,\mathcal{L}(\theta),\gamma)$
Algorithm 2 outlines the method for adjusting class proportions using the mixing parameters, $\alpha$, based on the computed class-wise losses. For each class, $i\in\{1,\ldots,k\}$, we first calculate the normalized loss $L_{i}$ by dividing the class-specific loss $\mathcal{L}_{i}(\theta)$ by the total cumulative loss summed over all classes. Each mixing parameter, $\alpha_{i}$, is then updated incrementally towards this normalized loss value $L_{i}$. The magnitude of the update step is controlled by the mixing rate, $\gamma$, determining how quickly the proportions adapt. Thus, classes exhibiting higher relative losses are progressively given greater emphasis in subsequent training epochs, ensuring a balanced reduction of errors across all classes.
@@
This cyclic selection procedure ensures that the required number of samples, $\alpha_{i}M$, for each class in every batch is maintained, even as $\alpha_{i}$ is adaptively updated across epochs.
@@
3 $\alpha_{i}\leftarrow\alpha_{i}+\gamma\left(L_{i}-\alpha_{i}\right)$ (Update mixing parameters)
@@
Algorithm 2 Updating Mixing Parameters Via Learn2Mix
@@ 4 Empirical Results
In this section, we present our empirical results on classification, regression, and image reconstruction tasks, across both benchmark and modified class imbalanced datasets. We first present the classification results on three benchmark datasets (MNIST (Deng, 2012), Fashion-MNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009)), and three standard datasets with manually imbalanced classes (Imagenette (Howard, 2020), CIFAR-100 (Krizhevsky et al., 2009), and IMDB (Maas et al., 2011)). We note that for the imbalanced case, we only introduce the manual class-imbalancing to the training dataset, $J$, where the test dataset, $\smash{K=\{(x_{j},y_{j})\}_{j=1}^{N_{\text{test}}}}$, is not changed. This choice ensures that the generalization performance of the network is benchmarked in a class-balanced setting. Next, for the regression task, we study two benchmark datasets with manually imbalanced classes (Wine Quality (Cortez et al., 2009), and California Housing (Géron, 2022)), and a synthetic mean estimation task, where the manual class-imbalancing parallels that of the classification case. Finally, we reconsider the MNIST, Fashion MNIST and CIFAR-10 datasets for image reconstruction, again with manual class-imbalancing. The comprehensive description of these datasets and class-imbalancing strategies is in Section C of the Appendix. For further performance verification, we include ablation studies on architecture, optimizer, batch size, learning rate, and worst-class error in Section B of the Appendix.
@@
The intuition behind the application of learn2mix to regression and reconstruction tasks stems from its ability to adaptively handle different data distributions. For regression tasks with a categorical variable taking $k$ distinct values, the samples from $J$ that correspond to each of the $k$ values, can be aggregated to obtain each class-specific training dataset, $J_{i}$. Here, each $J_{i}$ denotes a distinct underlying data distribution. As in the classification case, learn2mix will adaptively adjust the class-specific dataset proportions during training. For image reconstruction, we can similarly treat the $k$ distinct classes being reconstructed as the values taken by a categorical variable, paralleling the regression context. This formulation supports the adaptive adjustment of class proportions under learn2mix training.
@@
For the evaluations that follow, all training was performed on an NVIDIA GEForce RTX 3090 GPU. To ensure a fair comparison between learn2mix and classical training, we utilize the same learning rate, $\eta$, and neural network architecture with initialized parameters, $\theta$, across all experiments for a given dataset. Additionally, we train each neural network through learn2mix (with mixing rate $\gamma$) and classical training for $\smash{\underline{E}}$ seconds (or $E$ epochs), where $\smash{\underline{E}}$ is dataset dependent  [^1] Practically, we observe that choosing $\gamma\in[0.01,0.1]$ improves performance (see Section B of the Appendix).. In classification tasks, we also benchmark learn2mix and classical training versus ‘FCL training’, ‘SMOTE training’, ‘IS training’, and ‘CURR training’ (training using focal loss (Lin et al., 2017), SMOTE oversampling (Chawla et al., 2002), importance sampling (Katharopoulos and Fleuret, 2018), and curriculum learning (Hacohen and Weinshall, 2019) — see Sections D.3, D.4, D.5, and D.6 of the Appendix). The complete list of model architectures and hyperparameters is in Section D of the Appendix.
@@
Table 1: Test accuracies for learn2mix (L2M), classical (CL), FCL, SMOTE, IS, CURR training.
Elapsed Time: $\boldsymbol{\underline{t}=0.25\underline{E}}$ s
Dataset MNIST Fsh. MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Acc (L2M) $95.42\scriptstyle\pm 0.28$ $77.62\scriptstyle\pm 0.69$ $51.34\scriptstyle\pm 0.13$ $24.12\scriptstyle\pm 0.46$ $30.03\scriptstyle\pm 1.30$ $70.28\scriptstyle\pm 1.66$
Acc (CL) $93.14\scriptstyle\pm 0.47$ $74.13\scriptstyle\pm 0.73$ $49.26\scriptstyle\pm 0.15$ $15.55\scriptstyle\pm 0.13$ $23.23\scriptstyle\pm 1.99$ $50.13\scriptstyle\pm 0.15$
Acc (FCL) $91.32\scriptstyle\pm 0.57$ $74.08\scriptstyle\pm 0.75$ $47.90\scriptstyle\pm 0.22$ $20.11\scriptstyle\pm 0.37$ $27.15\scriptstyle\pm 1.13$ $50.30\scriptstyle\pm 0.67$
Acc (SMOTE) $92.41\scriptstyle\pm 0.71$ $73.67\scriptstyle\pm 0.61$ $47.76\scriptstyle\pm 0.15$ $23.19\scriptstyle\pm 0.46$ $23.93\scriptstyle\pm 2.35$ $50.94\scriptstyle\pm 0.05$
Acc (IS) $92.44\scriptstyle\pm 0.63$ $74.23\scriptstyle\pm 0.29$ $47.40\scriptstyle\pm 0.51$ $23.10\scriptstyle\pm 0.39$ $27.97\scriptstyle\pm 0.67$ $58.48\scriptstyle\pm 0.64$
Acc (CURR) $93.30\scriptstyle\pm 0.54$ $75.06\scriptstyle\pm 0.63$ $49.11\scriptstyle\pm 0.25$ $18.82\scriptstyle\pm 0.37$ $27.15\scriptstyle\pm 0.10$ $50.02\scriptstyle\pm 0.04$
Elapsed Time: $\boldsymbol{\underline{t}=0.5\underline{E}}$ s
Dataset MNIST Fsh. MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Acc (L2M) $97.61\scriptstyle\pm 0.15$ $83.16\scriptstyle\pm 0.87$ $55.84\scriptstyle\pm 0.19$ $33.64\scriptstyle\pm 0.42$ $46.80\scriptstyle\pm 0.54$ $76.02\scriptstyle\pm 2.77$
Acc (CL) $96.74\scriptstyle\pm 0.10$ $79.75\scriptstyle\pm 0.83$ $54.50\scriptstyle\pm 0.34$ $23.63\scriptstyle\pm 0.33$ $43.00\scriptstyle\pm 0.73$ $74.99\scriptstyle\pm 0.57$
Acc (FCL) $95.92\scriptstyle\pm 0.13$ $78.94\scriptstyle\pm 0.82$ $54.15\scriptstyle\pm 0.06$ $29.44\scriptstyle\pm 0.43$ $40.26\scriptstyle\pm 0.55$ $68.30\scriptstyle\pm 3.21$
Acc (SMOTE) $96.51\scriptstyle\pm 0.16$ $79.17\scriptstyle\pm 0.50$ $53.72\scriptstyle\pm 0.16$ $28.90\scriptstyle\pm 0.43$ $39.10\scriptstyle\pm 1.63$ $62.72\scriptstyle\pm 0.54$
Acc (IS) $96.60\scriptstyle\pm 0.25$ $79.65\scriptstyle\pm 0.38$ $52.56\scriptstyle\pm 0.39$ $28.52\scriptstyle\pm 0.32$ $42.61\scriptstyle\pm 2.61$ $74.00\scriptstyle\pm 0.81$
Acc (CURR) $96.53\scriptstyle\pm 0.16$ $79.08\scriptstyle\pm 0.58$ $53.49\scriptstyle\pm 0.33$ $27.26\scriptstyle\pm 0.79$ $39.48\scriptstyle\pm 2.22$ $71.68\scriptstyle\pm 0.55$
Elapsed Time: $\boldsymbol{\underline{t}=\underline{E}}$ s
Dataset MNIST Fsh. MNIST CIFAR-10 Imagenette CIFAR-100 IMDB
Acc (L2M) $98.46\scriptstyle\pm 0.14$ $85.85\scriptstyle\pm 0.47$ $60.49\scriptstyle\pm 0.26$ $42.95\scriptstyle\pm 0.33$ $54.50\scriptstyle\pm 0.73$ $82.38\scriptstyle\pm 0.59$
Acc (CL) $98.14\scriptstyle\pm 0.14$ $84.23\scriptstyle\pm 0.60$ $59.62\scriptstyle\pm 0.16$ $34.53\scriptstyle\pm 0.33$ $52.30\scriptstyle\pm 0.36$ $80.84\scriptstyle\pm 0.71$
Acc (FCL) $97.86\scriptstyle\pm 0.08$ $83.68\scriptstyle\pm 0.61$ $59.37\scriptstyle\pm 0.64$ $40.60\scriptstyle\pm 0.71$ $49.33\scriptstyle\pm 0.97$ $79.09\scriptstyle\pm 2.58$
Acc (SMOTE) $98.09\scriptstyle\pm 0.07$ $83.57\scriptstyle\pm 1.06$ $58.46\scriptstyle\pm 0.15$ $39.59\scriptstyle\pm 0.29$ $50.63\scriptstyle\pm 1.02$ $74.64\scriptstyle\pm 1.28$
Acc (IS) $98.14\scriptstyle\pm 0.14$ $84.33\scriptstyle\pm 0.29$ $57.44\scriptstyle\pm 0.42$ $35.33\scriptstyle\pm 0.43$ $52.83\scriptstyle\pm 0.34$ $79.08\scriptstyle\pm 0.57$
Acc (CURR) $98.13\scriptstyle\pm 0.05$ $83.32\scriptstyle\pm 0.43$ $59.15\scriptstyle\pm 0.42$ $35.26\scriptstyle\pm 0.48$ $50.88\scriptstyle\pm 0.79$ $80.04\scriptstyle\pm 0.25$
@@
Figure 2: Comparing model classification errors for learn2mix, classical, FCL, SMOTE, IS, and CURR training. The x-axis is the elapsed [training] time, while the y-axis is the classification error.
@@ 4.1 Classification Tasks
As illustrated in Table 1 and Figure 2, we observe a consistent trend across all considered classification benchmarks, whereby neural networks trained using learn2mix converge faster than their classically-trained, FCL-trained, SMOTE-trained, IS-trained, and CURR-trained counterparts. We first consider MNIST, and train LeNet-5 (Lecun et al., 1998) via the Adam optimizer (Kingma and Ba, 2014) and Cross Entropy Loss for $\underline{E}=50$ s, leveraging learn2mix, classical, FCL, SMOTE, IS, and CURR training. We see that the learn2mix-trained CNN converges faster, eclipsing a test accuracy of $98\%$ within $30$ s, whereas the remaining CNNs achieve this test accuracy after $40$ s. We next consider the more challenging Fashion MNIST dataset, and train Large LeNet-5 for $\underline{E}=50$ s with the Adam optimizer and Cross Entropy Loss, leveraging learn2mix, classical, FCL, SMOTE, IS, and CURR training. Paralleling MNIST, we observe that the learn2mix-trained CNN converges faster, yielding a test accuracy of $83\%$ within $20$ s, whereas the other CNNs achieve this test accuracy after $33$ s. The last class-balanced benchmark dataset we investigate is CIFAR-10, which offers a greater challenge than MNIST and Fashion MNIST. We train Large LeNet for $\underline{E}=200$ s using the Adam optimizer and Cross Entropy Loss, utilizing learn2mix, classical, FCL, SMOTE, IS, and CURR training. We observe that the learn2mix-trained CNN achieves faster convergence, yielding a test accuracy of $60\%$ after $170$ s, whereas the remaining CNNs exceed this test accuracy after $200$ s. Cumulatively, these evaluations demonstrate the efficacy of learn2mix training in settings with balanced classes, wherein the adaptive adjustment of class proportions accelerates convergence.
@@
We now consider several class-imbalanced training datasets. We first benchmark Imagenette, which comprises a subset of 10 classes from ImageNet (Deng et al., 2009), and modify the training dataset so the number of samples from each class, $i\in\{1,\ldots,k\}$, in $J$ decreases linearly. We train ResNet-18 (He et al., 2016) with the Adam optimizer and Cross Entropy Loss for $\underline{E}=230$ s, using learn2mix, classical, FCL, SMOTE, IS, and CURR training. We see the learn2mix-trained ResNet-18 converges faster, achieving a test accuracy of $40\%$ after $185$ s, whereas only the FCL-trained model achieves this test accuracy after $230$ s. We now consider CIFAR-100, and modify the training dataset so the number of samples from each class, $i\in\{1,\ldots,k\}$, in $J$ decreases logarithmically. We train MobileNet-V3 Small (Howard et al., 2019) for $\underline{E}=200$ s leveraging the Adam optimizer and Cross Entropy Loss, using learn2mix, classical, FCL, SMOTE, IS, and CURR training. We see that the learn2mix-trained MobileNet-V3 Small model converges faster, achieving a test accuracy of $50\%$ within $120$ s, whereas the other models exceed this test accuracy after $140$ s. As the $k=100$ mixing parameters are a small fraction of the total model parameters, this overhead is negligible. For IMDB, we modify the training dataset so the positive class keeps $30\%$ of its original samples. We train a transformer for $\underline{E}=150$ s with the Adam optimizer and Cross Entropy Loss, using learn2mix, classical, FCL, SMOTE, IS, and CURR training. We see the learn2mix-trained transformer converges faster, reaching a test accuracy of $70\%$ within $35$ s, whereas the other models exceed this test accuracy after $60$ s.
@@
In the above evaluations, we see learn2mix not only accelerates convergence, but also has a tighter alignment between training and test errors versus classical training. This correspondence indicates reduced overfitting, as learn2mix inherently adjusts class proportions based on class-specific error rates, $L_{i}$. By biasing the optimization procedure away from the original class distribution and towards $L_{i}$, learn2mix achieves improved generalization. We note this property is not unique to classification and also applies to regression and reconstruction (see Sections 4.2 and 4.3).
4.2 Regression Tasks
As illustrated in Table 2 and Figure 3, we observe that learn2mix maintains accelerated convergence in the regression context, wherein all the considered datasets are class imbalanced. We first consider the synthetic Mean Estimation dataset, which comprises sets of samples gathered from $k=4$ unique distributions and their associated means. Using the Adam optimizer and Mean Squared Error (MSE) Loss, we train a fully connected network for $E=500$ epochs on Mean Estimation via learn2mix and classical training. We see that the learn2mix-trained neural network converges rapidly, achieving a test error below $2.0$ after 100 epochs, at which point the classically-trained network has a test error of $13.0$. For the Wine Quality dataset, we modify the training dataset so the white wine class has $10\%$ of its original samples. Using the Adam optimizer and MSE Loss, we train a fully connected network for $E=300$ epochs on Wine Quality via learn2mix and classical training. We observe that the learn2mix-trained neural network yields faster convergence, achieving a test error below $2.5$ after 200 epochs, at which point the classically-trained network has a test error of $5.0$. Finally, on the California Housing dataset, we modify the training dataset such that three of the classes have $5\%$ of their original samples. Using the Adam optimizer and MSE Loss, we train a fully connected network for $E=1200$ epochs on California Housing via learn2mix and classical training. We again see that the learn2mix-trained network converges faster, achieving a test error below $0.8$ after 1200 epochs, while the classically-trained network has a test error of $0.99$. These empirical evaluations support our previous intuition pertaining to the extension of learn2mix to imbalanced regression settings.
4.3 Image Reconstruction Tasks
Per Table 2 and Figure 3, we note that the class-imbalanced image reconstruction tasks also observe faster convergence using learn2mix. For the MNIST case, we modify the training dataset such that half of the classes retain $20\%$ of their original samples. Leveraging the Adam optimizer and MSE Loss, we train an autoencoder for $E=40$ epochs on MNIST using learn2mix and classical training. We observe that the learn2mix-trained autoencoder exhibits improved convergence, achieving a test error less than $1.0$ after 35 epochs, which the classically-trained autoencoder achieves after 40 epochs. Correspondingly, for Fashion MNIST, we modify the training dataset such that half of the classes retain $20\%$ of their original samples (paralleling MNIST). Using the Adam optimizer and MSE Loss, we train an autoencoder for $E=70$ epochs on Fashion MNIST, leveraging learn2mix and classical training. We observe that the learn2mix-trained autoencoder converges faster, achieving a test error below $54.0$ after 50 epochs, which the classically-trained autoencoder achieves after 65 epochs. We also consider CIFAR-10, wherein we modify the training dataset such that all but two classes retain $20\%$ of their original samples. Utilizing the Adam optimizer and MSE Loss, we train an autoencoder for $E=110$ epochs on CIFAR-10, leveraging learn2mix and classical training. We observe that the learn2mix-trained autoencoder also converges faster and achieves a test error below $148.0$ after 100 epochs, which the classically-trained autoencoder achieves after 110 epochs.